problem:-

Case 1 - Sales = 40,000 units; Selling price per unit = $5.0; Fixed operating costs = $40,000.0; Unit variable cost = 40.0% of Selling price; Interest expense = $10,000.0; Taxes = 30.0%; No. of Common stock shares = 10,000.0 shares.

Case 2 - Sales = 60,000.0 units.

Case 3 - Sales = 20,000.0 units.

a) In case 1, what is your EBIT? EPS?

b) In case 3, what is the change in sales? EPS?

c) In case 2, what is the change in EBIT?

d) What is the DOL; DFL; DTL?
Answer #1

Solution:-

A. EBIT and EPS-

Home Insert Page Layout Formulas Data Review View Insert Function AutoSum Recently Financial Logical Text Date & Lookup & Use

B- Change in sales and in EPS-

Change in Sales = \frac{1,00,000 - 2,00,000}{2,00,000} * 100

Change in Sales = 50%

Hence, Decrease in sales = 50%.

Change in EPS = \frac{0.70 - 4.90}{4.90} * 100

Change in EPS = 85.71%

Decrease in EPS = 85.71%

C- Change in EBIT-

Change in EBIT = \frac{1,40,000 - 80,000}{80,000} * 100

Change in EBIT = 75%

Increase in EBIT is 75%.

D- To Calculate DOL, DFL, DTL-

Formula of Degree of operating Leverage = \frac{Contribution}{EBIT}

Formula of Degree of Financial Leverage = \frac{EBIT}{EBT}

Formula of Degree of Total Leverage = DOL * DFL

Calculation are done in above image.

If you have any query related to question then feel free to ask me in a comment. Thanks. Please rate.